# Delving into Temperature Scaling for Adaptive Conformal Prediction

### v2

@@ 5 Experiments
In this section, we first verify the effectiveness of ConfTS in both post-hoc and training conformal prediction methods across several benchmark datasets. Then, we investigate the adaptivity property of prediction sets employed with ConfTS, focusing on the SSCV performance and adaptiveness. In addition, we present ablation studies examining the impact of validation and conformal set sizes, as well as the effect of using different non-conformity scores to compute the efficiency gap in ConfTS.
@@ 5.1 Experimental setup
Datasets. In this work, we verify the effectiveness of ConfTS on CIFAR-100 (Krizhevsky et al., 2009), ImageNet (Deng et al., 2009), and ImageNet-V2 (Recht et al., 2019). On ImageNet, we split the test dataset, including 50,000 images, into 10,000 images for the calibration set and 40,000 images for the test set. On CIFAR-100 and ImageNet-V2, we split the test dataset, including 10,000 figures, into 4,000 figures for the calibration set and 6,000 for the test set. Additionally, we split the calibration set into two subsets of equal size: one subset is the validation set to optimize the temperature value with ConfTS, while the other half is the conformal set for conformal calibration.
Models. For ImageNet and ImageNet-V2, we employ 6 pre-trained classifiers from TorchVision (Paszke et al., 2019) – ResNet18, ResNet50, ResNet101 (He et al., 2016), DenseNet121 (Huang et al., 2017), VGG16 (Simonyan & Zisserman, 2015) and ViT-B-16 (Dosovitskiy et al., 2021). We also utilize the same model architectures for CIFAR-100 and train them from scratch. The models are trained for 100 epochs using SGD with a momentum of 0.9, a weight decay of 0.0005, and a batch size of 128. We set the initial learning rate as 0.1 and reduce it by a factor of 5 at 60 epochs. For conformal training, we set the smoothing parameter $T=0.1$, penalty term $\kappa=1$, and hyperparameter $\lambda=1$, using the same training setups. We conduct all the experiments on NVIDIA GeForce RTX 4090.
Conformal prediction algorithms. We leverage three adaptive conformal prediction methods: APS (Romano et al., 2020) and RAPS (Angelopoulos et al., 2021b) to generate prediction sets at error rate $\alpha\in\{0.1,0.05\}$. In addition, we set the regularization hyper-parameter for RAPS to be: $k_{reg}=1$ and $\lambda\in\{0.001,0.002,0.004,0.006,0.01,0.015,0.02\}$. For the evaluation metrics, we employ coverage, average size, and SSCV (see Appendix A) to assess the performance of prediction sets. All experiments are repeated 20 times with different seeds, and we report average performances.
Table 2: Performance of ConfTS using APS and RAPS on ImageNet dataset. We repeat each experiment for 20 times. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that our ConfTS can improve the performance of APS and RAPS, maintaining the desired coverage rate.
Model Score $\alpha=0.1$ $\alpha=0.05$
Coverage Average $\downarrow$ Coverage Average $\downarrow$
@@
ResNet18 APS 0.900 / 0.900 14.09 / 7.531* 0.951 / 0.952 29.58 / 19.59*
RAPS 0.900 / 0.900 9.605 / 5.003* 0 .950 / 0.950 14.72 / 11.08*
ResNet50 APS 0.899 / 0.900 9.062 / 4.791* 0.950 / 0.951 20.03 / 12.22*
RAPS 0.899 / 0.900 5.992 / 3.561* 0.950 / 0.951 9.423 / 5.517*
ResNet101 APS 0.900 / 0.899 6.947 / 4.328* 0.950 / 0.950 15.73 / 10.51*
RAPS 0.900 / 0.899 4.819 / 3.289* 0.950 / 0.950 7.523 / 5.091*
DenseNet121 APS 0.900 / 0.899 9.271 / 4.746* 0.950 / 0.949 20.37 / 11.47*
RAPS 0.900 / 0.900 6.602 / 3.667* 0.949 / 0.949 10.39 / 6.203*
VGG16 APS 0.901 / 0.901 11.73 / 6.057* 0.951 / 0.951 23.71 / 14.78*
RAPS 0.901 / 0.900 8.118 / 4.314* 0.950 / 0.950 12.27 / 8.350*
ViT-B-16 APS 0.900 / 0.901 14.64 / 2.315* 0.951 / 0.950 36.72 / 9.050*
RAPS 0.902 / 0.901 6.889 / 1.800* 0.950 / 0.950 12.63 / 3.281*
Average APS 0.900 / 0.900 10.96 / 4.961* 0.950 / 0.950 24.36 / 12.94*
RAPS 0.900 / 0.900 7.000 / 3.606* 0.950 / 0.950 11.16 / 6.587*
@@ 5.2 Main results
ConfTS improves current adaptive conformal prediction methods.
@@
In Table 2, we present the performance of APS and RAPS ($\lambda=0.001$) with ConfTS on the ImageNet dataset. A salient observation is that ConfTS drastically improves the efficiency of adaptive conformal prediction, while maintaining the marginal coverage. For example, on the ViT model at $\alpha=0.05$, ConfTS reduces the average size of APS by 7 times - from 36.72 to 5.759. Averaged across six models, ConfTS improves the efficiency of APS by 58.3% at $\alpha=0.1$. We observe similar results on CIFAR-100 and ImageNet-V2 dataset in Appendix G and Appendix H. Moreover, our ConfTS remains effective for RAPS across various penalty terms on ImageNet as shown in Appendix I. Furthermore, in Appendix J, we demonstrate that ConfTS can lead to small prediction sets for SAPS (Huang et al., 2024) – another adaptive conformal prediction technique. Overall, empirical results show that ConfTS consistently improves the efficiency of existing adaptive conformal prediction methods across various networks.
Figure 2: The SSCV performance of ConfTS using APS on ImageNet dataset. A smaller SSCV is better. The results show that ConfTS can improve the conditional coverage performance of APS.
Table 3: The performance of ConfTS when applied to models trained with ConfTr loss on CIFAR-100. We repeat each experiment for 20 times. “*” denotes significant improvement (two-sample t-test at a 0.1 confidence level). Bold numbers are superior results. $\downarrow$” indicates smaller values are better. The results show that ConfTS boosts ConfTr: it generates efficient prediction sets for APS and RAPS.
Model Score $\alpha=0.1$ $\alpha=0.05$
Coverage Average size $\downarrow$ Coverage Average size $\downarrow$
Baseline / ConfTS
ResNet18 APS 0.899 / 0.899 6.670 / 5.827* 0.949 / 0.949 11.49 / 10.35*
RAPS 0.900 / 0.900 5.848 / 4.799* 0.951 / 0.951 8.431 / 8.189*
ResNet50 APS 0.900 / 0.901 5.754 / 5.075* 0.951 / 0.951 9.911 / 9.150*
RAPS 0.900 / 0.900 5.186 / 4.324* 0.951 / 0.949 7.868 / 7.507*
ConfTS boosts training-time conformal prediction method.
@@
Previous work (Stutz et al., 2022) proposes Conformal Training (ConfTr) which enhances the efficiency of prediction sets during training process. In this part, we investigate how our ConfTS interacts with ConfTr. In particular, we leverage ResNet18 and ResNet50 trained with ConfTr loss for 100 epochs on CIFAR-100, generating prediction sets with APS and RAPS ($\lambda=0.001$) method at error rate $\alpha\in\{0.1,0.05\}$. In Table 3, we present the performance of ConfTS when applied to networks trained with ConfTr loss. The results show that our ConfTS can boost the performance of ConfTr. For instance, on the ResNet50 model trained with ConfTr loss, at an error rate $\alpha=0.1$, ConfTS reduces the average size of RAPS from 5.186 to 4.324. In summary, our findings highlight that ConfTS can effectively improve both post-hoc and training methods of conformal prediction.
ConfTS enhances the conditional coverage.
The conditional coverage (Vovk, 2012) requires conformal prediction methods to satisfy the marginal coverage at instance level. The size-stratified coverage violation (SSCV) (Angelopoulos et al., 2021b) is often employed to evaluate the conditional coverage of prediction sets:
$\text{SSCV}=\sup_{j}|\frac{\left|\{i\in S_{j}:y_{i}\in\mathcal{C}\left(\bm{x}_{i}\right)\}\right|}{\left|S_{j}\right|}-(1-\alpha)|,$
where $\{S_{i}\}_{i=1}^{N_{s}}$ is a disjoint set-size strata, satisfying $\bigcup_{i=1}^{N_{s}}S_{i}=\{1,2,\cdots,|\mathcal{Y}|\}$. Specifically, a lower SSCV value indicates better conditional coverage performance. Following prior work (Angelopoulos et al., 2021b), we set the partitioning of the set sizes as: 0-1, 2-3, 4-10, 11-100, and 101-1000. Figure 2 presents the SSCV performance of ConfTS using APS on ImageNet at $\alpha=0.05$. The results show that ConfTS can enhance conditional coverage in most cases. For example, on ResNet50, ConfTS reduces the SSCV of APS from 0.033 to 0.025. Overall, these findings highlight that our ConfTS method improves both conditional coverage and the efficiency of adaptive conformal prediction.
ConfTS maintains the adaptiveness.
Adaptiveness (Romano et al., 2020; Angelopoulos et al., 2021b; Seedat et al., 2023) requires prediction sets to communicate instance-wise uncertainty: easy examples should obtain smaller sets than hard ones. In this part, we examine the impact of ConfTS on the adaptiveness of prediction sets and measure the instance difficulty by the order of the ground truth $o(y,\pi(\bm{x}))$. Specifically, we partition the sample by label order: 1, 2-3, 4-6, 7-10, 11-100, 101-1000, following (Angelopoulos et al., 2021b). Figure 3(a) and Figure 3(b) present the adaptiveness performance of ConfTS with APS score, using ResNet18 and ResNet50 on the ImageNet at $\alpha=0.1$. A salient observation is that prediction sets, when applied with ConfTS, satisfy the adaptiveness property. Notably, employing ConfTS can promote smaller prediction sets for all examples ranging from easy to hard. Overall, the results demonstrate that APS with ConfTS succeeds in producing adaptive prediction sets: examples with lower difficulty obtain smaller prediction sets on average.
Ablation study on the size of validation and calibration set.
In the experiment, ConfTS splits the calibration data into two subsets: validation set for tuning the temperature and conformal set for conformal calibration. In this part, we analyze the impact of this split on the performance of ConfTS by varying the validation and conformal dataset sizes from 3,000 to 8,000 samples while maintaining the other part at 5,000 samples. We use ResNet18 and ResNet50 on ImageNet, with APS at $\alpha=0.1$. Figure 3(c) and 3(d) show that the performance of ConfTS remains consistent across different conformal dataset sizes and validation dataset sizes. Based on these results, we choose a calibration set including 10000 samples and split it into two equal subsets for the validation and conformal set. In summary, the performance of ConfTS is robust to variations in the validation dataset and conformal dataset size.
(a) ResNet18
(b) ResNet50
Figure 3: (a)&(b): Average sizes of examples with different difficulties using APS on ResNet18 and ResNet50 respectively. Results show that ConfTS can maintain adaptiveness. (b)&(c) Average sizes of APS employed with ConfTS under various sizes of (b) conformal dataset (c) validation dataset. Results show that our ConfTS is robust to variations in the validation and conformal dataset size.
@@
Table 4: The performance of ConfTS using various non-conformity scores to compute the efficiency gap. We consider standard APS and RAPS score as well as their non-randomized variants. Each experiment is repeated 20 times. “Avg.size” and “Cov.” represent the results of average size and coverage, and ’Base’ presents the results without ConfTS. $\downarrow$” indicates smaller values are better. “” and “” indicate the performance is superior/inferior to the baseline. Bold numbers are superior results. Results show that using the non-randomized APS score achieves the overall best performance.
@@
Ablation study on the non-conformity score in ConfTS.
In this ablation, we compare the performance of ConfTS trained with various non-conformity scores in Eq. (8), including standard APS and RAPS, as well as their non-randomized variants. Table 4 presents the performance of prediction sets generated by standard APS and RAPS ($\lambda=0.001$) methods with different variants of ConfTS, employing ResNet18 and ResNet50 on ImageNet. The results show that ConfTS with randomized scores fails to produce efficient prediction sets, while non-randomized scores result in small prediction sets. This is because the inclusion of the random variable $u$ leads to the wrong estimation of the efficiency gap, thereby posing challenges to the optimization process in ConfTS. Moreover, randomized APS consistently performs better than randomized RAPS, even in the case of using the standard RAPS to generate prediction sets. Overall, our findings show that ConfTS with the non-randomized APS outperforms the other scores in enhancing the efficiency of prediction sets.
6 Related Work
Conformal prediction. Conformal prediction (Papadopoulos et al., 2002; Vovk et al., 2005) is a statistical framework for uncertainty qualification. Previous works have applied conformal prediction across various domains, including regression (Lei & Wasserman, 2014; Romano et al., 2019), image classification (Sadinle et al., 2019; Angelopoulos et al., 2021b; Huang et al., 2024), hyperspectral image classification (Liu et al., 2024), object detection (Angelopoulos et al., 2021a; Teng et al., ), and large language models (Kumar et al., 2023). Conformal prediction is also deployed in other real-world applications, such as human-in-the-loop decision-making (Straitouri et al., 2023; Cresswell et al., 2024), automated vehicles (Bang et al., 2024), and scientific computation (Moya et al., 2024).
Some methods leverage post-hoc techniques to enhance the performance of prediction sets (Romano et al., 2020; Angelopoulos et al., 2021b; Ghosh et al., 2023; Huang et al., 2024). For example, Adaptive Prediction Sets (APS) (Romano et al., 2020) calculates the score by accumulating the sorted softmax values in descending order. However, the softmax probabilities typically exhibit a long-tailed distribution, and thus those tail classes are often included in the prediction sets. To alleviate this issue, Regularized Adaptive Prediction Sets (RAPS) (Angelopoulos et al., 2021b) exclude tail classes by appending a penalty to these classes, resulting in efficient prediction sets. Moreover, these post-hoc methods often employ temperature scaling for better calibration performance (Angelopoulos et al., 2021b; Lu et al., 2022; Gibbs et al., 2023; Lu et al., 2023). In our work, we show that the bulk of confidence calibration methods increase the average size of the prediction set, which motivates us to design a variant of temperature scaling, i.e., ConfTS, to enhance the efficiency of prediction sets.
Some works propose training time regularizations to improve the efficiency of conformal prediction (Colombo & Vovk, 2020; Stutz et al., 2022; Einbinder et al., 2022; Bai et al., ; Correia et al., 2024). For example, uncertainty-aware conformal loss function (Einbinder et al., 2022) optimizes the efficiency of prediction sets by encouraging the non-conformity scores to follow a uniform distribution. Moreover, conformal training (Stutz et al., 2022) constructs efficient prediction sets by prompting the threshold to be less than the non-conformity scores. In addition, information-based conformal training (Correia et al., 2024) incorporates side information into the construction of prediction sets. In this work, we mainly focus on enhancing adaptive conformal prediction in a post-hoc manner, which is parameter-efficient and requires low computational resources. Notably, our findings show that ConfTS can effectively improve the performance of both post-hoc and training-time conformal prediction methods.
Confidence calibration. Confidence calibration has been studied in various contexts in recent years. Some works address the miscalibration problem by post-hoc methods, including histogram binning (Zadrozny & Elkan, 2001) and Platt scaling (Platt et al., 1999). Besides, regularization methods like entropy regularization (Pereyra et al., 2017) and focal loss (Mukhoti et al., 2020) are also proposed to improve the calibration performance of deep neural networks. The most related work explored the influence of confidence calibration on confidence intervals in binary classification settings (Gupta et al., 2020). Our work aligns with and extends these findings by focusing on multi-class classification scenarios, which are more prevalent in practical applications. Moreover, we provide a detailed examination of how the temperature value affects the performance of adaptive conformal prediction.

### v3

@@ 4.1 Experimental setup
Datasets. We evaluate ConfTS on both image and text classification tasks. For image classification, we employ CIFAR-100 (Krizhevsky et al., 2009), ImageNet (Deng et al., 2009), and ImageNet-V2 (Recht et al., 2019). For text classification, we utilize AG news (Zhang et al., 2015) and DBpedia (Auer et al., 2007) datasets. For ImageNet, we split the test dataset containing 50,000 images into 10,000 images for calibration and 40,000 for testing. For CIFAR-100 and ImageNet-V2, we split their test datasets, each containing 10,000 figures, into 4,000 figures for calibration and 6,000 for testing. For text datasets, we split each test dataset equally between calibration and testing. Additionally, we split the calibration set into two subsets of equal size: one subset is the validation set to optimize the temperature value with ConfTS, while the other half is the conformal set for conformal calibration.
Models. For ImageNet and ImageNet-V2, we employ 6 pre-trained classifiers from TorchVision (Paszke et al., 2019) – ResNet18, ResNet50, ResNet101 (He et al., 2016), DenseNet121 (Huang et al., 2017), VGG16 (Simonyan & Zisserman, 2015) and ViT-B-16 (Dosovitskiy et al., 2021). We also utilize the same model architectures for CIFAR-100 and train them from scratch. For text classification, we finetune a pre-trained BERT (Devlin, 2018) and GPT-Neo-1.3B (Black et al., 2021) on each dataset. The model architecture consists of a frozen pre-trained encoder followed by a trainable linear classifier layer. For each dataset, we employ the AdamW optimizer with a learning rate of 2e-5. The training is conducted over 3 epochs with a batch size of 32. The models are trained for 100 epochs using SGD with a momentum of 0.9, a weight decay of 0.0005, and a batch size of 128. We set the initial learning rate as 0.1 and reduce it by a factor of 5 at 60 epochs.
Conformal prediction algorithms. We leverage three adaptive conformal prediction methods, APS and RAPS, to generate prediction sets at error rate $\alpha\in\{0.1,0.05\}$. In addition, we set the regularization hyper-parameter for RAPS to be: $k_{reg}=1$ and $\lambda\in\{0.001,0.002,0.004,0.006,0.01,0.015,0.02\}$. For the evaluation metrics, we employ coverage and average size to assess the performance of prediction sets. All experiments are repeated 20 times with different seeds, and we report average performances.
Table 2: Performance of ConfTS using APS and RAPS on ImageNet dataset. The tuned $T$ is the temperature value optimized by our loss function. We repeat each experiment for 20 times. $\downarrow$” indicates smaller values are better. Bold numbers are superior results. Results show that our ConfTS can improve the performance of APS and RAPS, maintaining the desired coverage rate.
Model Error rate Tuned $T$ APS RAPS
Coverage Average size $\downarrow$ Coverage Average size $\downarrow$
@@
ResNet18 $\alpha=0.1$ 0.593 0.900 / 0.900 14.09 / 7.531 0.900 / 0.900 9.605 / 5.003
$\alpha=0.05$ 0.591 0.951 / 0.952 29.58 / 19.59 0.950 / 0.950 14.72 / 11.08
ResNet50 $\alpha=0.1$ 0.705 0.899 / 0.900 9.062 / 4.791 0.899 / 0.900 5.992 / 3.561
$\alpha=0.05$ 0.709 0.950 / 0.951 20.03 / 12.22 0.950 / 0.951 9.423 / 5.517
ResNet101 $\alpha=0.1$ 0.793 0.900 / 0.899 6.947 / 4.328 0.900 / 0.899 4.819 / 3.289
$\alpha=0.05$ 0.785 0.950 / 0.950 15.73 / 10.51 0.950 / 0.950 7.523 / 5.091
DenseNet121 $\alpha=0.1$ 0.659 0.900 / 0.899 9.271 / 4.746 0.900 / 0.900 6.602 / 3.667
$\alpha=0.05$ 0.675 0.950 / 0.949 20.37 / 11.47 0.949 / 0.949 10.39 / 6.203
VGG16 $\alpha=0.1$ 0.604 0.901 / 0.901 11.73 / 6.057 0.901 / 0.900 8.118 / 4.314
$\alpha=0.05$ 0.627 0.951 / 0.951 23.71 / 14.78 0.950 / 0.950 12.27 / 8.350
ViT-B-16 $\alpha=0.1$ 0.517 0.900 / 0.901 14.64 / 2.315 0.902 / 0.901 6.889 / 1.800
$\alpha=0.05$ 0.482 0.951 / 0.950 36.72 / 9.050 0.950 / 0.950 12.63 / 3.281
@@ 4.2 Main results
ConfTS improves current adaptive conformal prediction methods. In Table 2, we present the performance of APS and RAPS ($\lambda=0.001$) with ConfTS on the ImageNet dataset. A salient observation is that ConfTS drastically improves the efficiency of adaptive conformal prediction, while maintaining the marginal coverage. For example, on the ViT model at $\alpha=0.05$, ConfTS reduces the average size of APS by 7 times - from 36.72 to 5.759. Averaged across six models, ConfTS improves the efficiency of APS by 58.3% at $\alpha=0.1$. We observe similar results on CIFAR-100 and ImageNet-V2 dataset in Appendix I and Appendix H. Moreover, our ConfTS remains effective for RAPS across various penalty terms on ImageNet as shown in Appendix J. Furthermore, in Appendix K, we demonstrate that ConfTS can lead to small prediction sets for SAPS (Huang et al., 2024), a recent technique of adaptive conformal prediction. In addition, we find that the tuned temperature values are generally smaller than 1.0 and different for various settings, which demonstrates the importance of the automatic method. Overall, empirical results show that ConfTS consistently improves the efficiency of existing adaptive conformal prediction methods.
Table 3: The average size of APS and RAPS using ConfPS and ConfVS. ConfPS and ConfVS are the variants of Platt scaling and vector scaling, optimized by our ConfTS loss. “” and “” indicate the performance is superior/inferior to the baseline. The results show that by rescaling the logits with ConfPS and ConfVS, the algorithm can construct efficient prediction sets, demonstrating the generality of our loss function.
Dataset Model APS RAPS
Baseline ConfTS ConfPS ConfVS Baseline ConfTS ConfPS ConfVS
ImageNet ResNet50 9.062 4.791 2.571 4.564 5.992 3.561 2.446 3.303
DenseNet121 9.271 4.746 3.169 5.345 6.602 3.667 3.224 3.683
VGG16 11.73 6.057 3.729 7.020 8.118 4.314 3.558 4.745
ViT-B-32 14.64 2.315 1.743 4.797 6.899 1.800 1.575 2.549
AG news BERT 2.105 1.886 1.808 1.979 2.004 1.802 1.794 1.949
GPT-Neo-1.3B 2.022 1.911 1.749 1.897 2.018 1.909 1.728 1.884
Dbpedia BERT 3.557 2.905 2.96 3.869 3.458 2.908 2.837 3.744
GPT-Neo-1.3B 3.171 2.178 1.826 1.884 3.137 2.144 1.768 2.415
Average 13.890 6.697 4.889 7.839 9.557 5.526 4.733 6.068
Table 4: The average size of APS and RAPS with various post-hoc calibration methods optimized by our loss and ConfTr loss. “” and “” indicate the performance is superior/inferior to the baseline. Bold numbers are superior results between two loss functions. The results show that ConfTS loss achieves better performance than ConfTr loss in most cases.
Model Baseline ConfTS ConfPS ConfVS
Our loss ConfTr loss Our loss ConfTr loss Our loss ConfTr loss
ResNet50 APS 9.062 4.719 / 8.864 2.571 / 2.657 4.564 / 4.471
RAPS 5.992 3.561 / 5.980 2.446 / 2.500 3.303 / 3.333
VGG16 APS 11.73 6.057 / 9.822 3.729 / 4.193 7.020 / 6.757
RAPS 8.118 4.314 / 6.825 3.558 / 3.921 4.745 / 4.742
Average 8.726 4.663 / 7.873 3.076 / 3.318 4.908 / 4.826
Our method can work with other post-hoc calibration methods. We extend the application of ConfTS loss (Eq. (8)) to other post-hoc calibration methods. We introduce conformal Platt scaling (dubbed ConfPS) and conformal vector scaling (dubbed ConfVS), where the parameters are optimized using ConfTS loss. We employ ResNet50 and VGG16 models on the ImageNet dataset for image task, as well as BERT and GPT-Neo-13B on the DBpedia dataset for text task. The error rate is $\alpha=0.1$. Table 3 shows that both ConfPS and ConfVS can help construct efficient prediction sets. This indicates that replacing cross-entropy loss with ConfTS loss in post-hoc calibration methods consistently enhances the efficiency of adaptive conformal prediction. Overall, these results validate the effectiveness of ConfTS loss across different calibration methods.
@@
Our loss function outperforms ConfTr loss. Previous work (Stutz et al., 2022) proposes Conformal Training (ConfTr), which enhances prediction set efficiency during training through a novel ConfTr loss function. In this part, We compare the performance of ConfTS, ConfPS, and ConfVS when trained with both ConfTr loss and our proposed ConfTS loss. Using ResNet50 and VGG16 models on ImageNet, we generate prediction sets with APS and RAPS at an error rate $\alpha=0.1$. The results in Table 4 demonstrate that while both loss functions improve prediction set efficiency, our loss normally achieves better performance than ConfTr loss. For example, with the ResNet50 model, ConfTS loss reduces the average size of APS to 4.791, compared to 8.864 when using ConfTr loss. Overall, the proposed loss function is superior to the ConfTr loss in optimizing the calibration methods.
@@
Table 5: The performance of ConfTS using various non-conformity scores to compute the efficiency gap. We consider standard APS and RAPS score as well as their non-randomized variants. Each experiment is repeated 20 times. “Avg.size” and “Cov.” represent the results of average size and coverage, and ’Base’ presents the results without ConfTS. $\downarrow$” indicates smaller values are better. “” and “” indicate the performance is superior/inferior to the baseline. Bold numbers are superior results.
@@
Ablation study on the non-conformity score in ConfTS. In this ablation, we compare the performance of ConfTS trained with various non-conformity scores in Eq. (8), including standard APS and RAPS, as well as their non-randomized variants. Table 5 presents the performance of prediction sets generated by standard APS and RAPS ($\lambda=0.001$) methods with different variants of ConfTS, employing ResNet18 and ResNet50 on ImageNet. The results show that ConfTS with randomized scores fails to produce efficient prediction sets, while non-randomized scores result in small prediction sets. This is because the inclusion of the random variable $u$ leads to the wrong estimation of the efficiency gap, thereby posing challenges to the optimization process in ConfTS. Moreover, randomized APS consistently performs better than randomized RAPS, even in the case of using the standard RAPS to generate prediction sets. Overall, our findings show that ConfTS with the non-randomized APS outperforms the other scores in enhancing the efficiency of prediction sets.